Irving has 35 coins, consisting of only dimes and nickels. If the value of his coins is $3.30, how many nickels and dimes does he have?
Answer by mananth(16946)

let dimes be x
nickels will be 35-x
..
Total value = $3.30 = 330 cents
10x+5(35-x)=330
10x+175-5x=330
5x=330-175
5x=155
x= 155/5
x=31 number of dimes
nickels will be 2